OpenStudy (anonymous):

Chico walks 3/5 mile to the library. Nikki walks 1/4 the distance that Chico walks to the library. How far does Nikki walk?

OpenStudy (anonymous):

can u plz help pooja

OpenStudy (anonymous):

hint: multiply the fractions

pooja195 (pooja195):

\[\huge~\rm~\bf~\frac{ 3 }{5}\times \frac{ 1 }{ 4}=? \]

OpenStudy (anonymous):

If Chico walks 3/5 of the way, he is 60% away when you convert it into a percentage. Then you have to find 1/4 of that or 25% of 60%.

OpenStudy (anonymous):

3/20

pooja195 (pooja195):

good

OpenStudy (anonymous):

thanku